### v2

# Human as Points: Explicit Point-based 3D Human Reconstruction from Single-view RGB Images

Abstract
[…]
Index Terms:
Human reconstruction, single-view image, point cloud, depth, SMPL.
I Introduction
Reconstructing the 3D geometric structures of human bodies from single-view 2D RGB images has been a fundamental and long-standing task, which lays the foundation for numerous downstream applications, such as immersive communication, video games, and virtual/augmented reality. However, it still remains highly challenging to produce accurate and detailed reconstruction results, due to the difficulty in inferring complete 3D human geometry information from heavily occluded regions in single-view image capture.
In recent years, motivated by the efficiency and flexibility of deep implicit functions [11, 37, 42] to represent unconstrained-topology 3D shapes with fine-level geometric details, pixel-aligned implicit reconstruction pipelines [50, 51] have become the dominating processing paradigm due to their ability to recover high-fidelity geometry for clothed humans. However, these approaches typically overfit the limited training data due to the lack of human body priors, resulting in degenerate body structures and broken limbs when encountering novel poses or clothing styles. To overcome such limitations, the more recent trend [68, 65] is to constrain the fully implicit learning mechanism by incorporating parametric 3D human body models [33, 45] as the explicit shape priors. However, although combining implicit representations with explicit geometric constraints is considered as a more promising exploration direction, it also poses additional challenges and higher requirements for the design of the specific combination scheme. The fusion of learned features or intermediate results inferred from input image pixels and the estimated parametric model typically leads to geometric over-smoothing and thus weakens surface details. More importantly, the topological inconsistency between the chosen parametric template model and the target human shape (especially with loose clothing or complex accessories) usually causes over-constraining effects. The latest effort [64] jointly exploits the representation capability of implicit fields and robustness of explicit human body models, but the quality of the resulting reconstruction can still be constrained by unsatisfactory rectifications of parametric body models [45] and inaccurate predictions of normal maps.
In practice, there are several aspects of key consideration when evaluating the performance of the human reconstruction pipeline:
  1. 1.
    flexibility to support unconstrained-topology modeling of human body shapes with arbitrary clothing;
  2. 2.
    generalizability to unseen data distributions (e.g., novel poses and clothes);
  3. 3.
    robustness to avoid non-human reconstruction results (e.g., unnatural poses, degenerate body structures, and broken limbs);
  4. 4.
    capability to represent and capture fine-level details of surface geometry in 3D space.
In this paper, we seek to develop a learning-based single-view human reconstruction framework that simultaneously meets the above critical requirements, making it a more promising processing paradigm with greater potential compared with previous methods. Architecturally, we construct HaP, a fully explicit point-based modeling pipeline featuring direct point cloud estimation (from the 2D image), manipulation, generation, and refinement in the explicit 3D geometric space.
More specifically, given an input RGB image, our process begins with depth estimation to infer a partial 3D point cloud representing high-fidelity visible-view geometry. Meanwhile, we also estimate the SMPL model from the 2D image space to provide the missing human body information. Under the assumption that the estimated depth maps are generally accurate, the preceding depth-inferred partial 3D points are further exploited for pose correction through a specialized SMPL rectification procedure to make the estimated SMPL model register well with the points in the 3D space. Directly merging the 3D information of both the depth-inferred partial points and the estimated SMPL model can coarsely form a complete 3D human point cloud with rich texture and wrinkle details; however, there are still apparent gaps between the partial points and the estimated SMPL model, which makes the combination of the partial points and the estimated SMPL model unnatural and possibly problematic. To this end, we customize a diffusion-style point cloud generation framework to learn the latent spatial distribution of realistic human bodies based on the coarse merging 3D human points, to generate 3D human bodies with clothing and pose information consistent with the depth-inferred partial 3D points and the SMPL model. Moreover, we propose a refinement stage, including learning-based displacement estimation and a simple yet effective depth replacement strategy, to further enhance the geometric quality. Finally, a high-quality mesh model can be derived from the point cloud using standard surface reconstruction methods, such as screened Poisson surface reconstruction [25].
[…]
With the design of the dedicated components, the proposed HaP framework comprehensively addresses the concerns mentioned above. Concretely,
  1. 1.
    we utilize point clouds as the intermediate representation, which is known to be flexible [47, 48] for modeling arbitrary-topology geometric structures;
  2. 2.
    the process of inferring coarse 3D geometry information from the 2D image consists of 2D depth estimation and SMPL estimation. The former, as a richly investigated and relatively mature task, shows satisfactory generalizability. The latter improves reconstruction robustness by injecting human body priors without destroying clothing cues mined from the specific input image; and
  3. 3.
    it is more straightforward and effective to rectify poses and capture surface details when the point-based learning process is directly implemented in the original 3D geometric space.
Extensive experiments demonstrate that our HaP outperforms state-of-the-art approaches both quantitatively and qualitatively (see Fig. 1).
In summary, the major contributions of this paper are as follows.
  • we propose a new pipeline, HaP, to explicitly generate human bodies in point clouds through a diffusion process conditioned on depth maps and SMPL models in 3D space;
  • we propose an effective SMPL rectification module to optimize SMPL models to accurate poses;
  • we contribute a new human scan dataset comprising fine-grained, detailed scans to facilitate future research endeavors, as illustrated in Fig. 2.
The remainder of this paper is organized as follows. Section II provides a comprehensive review of the existing literature, including monocular depth estimation, point cloud generation, human pose and shape estimation and single-view human reconstruction. Section III introduces our proposed HaP in detail. In Section IV, we conduct extensive experiments and ablation studies to evaluate the effectiveness of HaP. Finally, Section V concludes this paper.
II Related Work
The HaP pipeline consists of three main modules: depth estimation from single-view RGB images, SMPL estimation and rectification using single-view RGB images, and conditional human point cloud generation. Based on these modules, we provide a concise overview of the related work in Monocular Depth Estimation, SMPL Estimation and Rectification, and Point Cloud Representation. Additionally, we briefly review other methods in Clothed Human Reconstruction.
II-A Monocular Depth Estimation
Predicting depth maps from single-view RGB images is a challenging task. Over the years, various approaches [44, 52, 13, 28, 43, 62] have been developed to tackle this problem. Make3D [52] divides the image into small homogeneous patches and uses a Markov random field to infer the plane parameters, which capture the location and orientation of each patch in 3D space. Eigen et al[13] proposed a pioneering deep learning-based method to learn an end-to-end mapping from RGB images to depth maps. BTS [28] designs novel local planar guidance layers for more effectively utilizing the densely encoded feature at multiple stages when decoding. P3Depth [43] iteratively and selectively leverages the information from coplanar pixels to improve the quality of the predicted depth. Xie et al[62] found that Masked image modeling (MIM) could also achieve state-of-the-art performance on depth estimation.
Particularly, various methods have been proposed for estimating human depth maps from single-view images. Tang et al. [56] employed a segmentation net and a skeleton net to generate human body parts and keypoints heatmaps. These heatmaps were used to compute the base shape and detail shape via a depth net. Tan et al. [55] proposed a self-supervised method that utilizes the predicted SMPL models from videos. Similarly, HDNet [24] predicts the corresponding dense pose of each image. Both Tan et al. [55] and Jafarian et al. [24] warped body parts to different frames of a video and supervised the depth maps with a photo-consistency loss, as the depth map would not change much among different frames.
Building on the latest monocular depth estimation method [62], we extract precise 3D information from the front view of the human body. This depth information acts as supervision for SMPL rectification and provides essential conditions for the diffusion model, significantly enhancing the accuracy of the overall reconstruction. By emphasizing depth as a cornerstone of the task, this approach achieves substantial improvements over previous methodologies.
II-B SMPL Estimation and Rectification
The human body structure provides essential prior information for human body reconstruction. Estimating 2D keypoints [15, 8] or human parsing [10, 29] can offer valuable insight into human structure. However, both 2D keypoints and human parsing lack the spatial information necessary for precise 3D reconstruction. In contrast, predicting parametric models, such as SMPL(-X) [33], from RGB images enables a more comprehensive understanding of human pose and shape, offering stronger prior knowledge for 3D human reconstruction.
Zhang et al.[67] proposed a regression-based approach that uses a feature pyramid and corrects the predicted parameters by aligning the SMPL mesh and image. PIXIE [16] introduces a moderator that merges the body, face, and hand features of experts according to their confidence weights. Kocabas et al.[27] proposed an adversarial learning framework to produce kinematical SMPL motion sequences in the absence of in-the-wild 3D ground truth. METRO [30] attempts to model the interactions of vertex-vertex pairs and vertex-joint connections, and finally outputs human joint coordinates and SMPL vertices together.
However, predicted SMPL(-X) [33, 45] models sometimes exhibit misalignment issues when observed in 3D space. Xiu et al[64, 65] utilized the rendering-based optimization method to rectify the SMPL poses. Zheng et al. [68] optimized the SMPL parameters in the 3D implicit space. In this paper, we propose a novel SMPL rectification module that minimizes the distance between the visible part of the SMPL model and the depth-inferred partial point cloud in 3D space, and generate SMPL models that are more accurate and better-aligned with the given human body.
II-C Point Cloud Representation
Point clouds are an efficient and effective representation of 3D shapes, often serving as a precursor for generating other 3D formats, such as meshes and voxels. They have wide-ranging applications across various fields.
Point Cloud Generation. Generating point clouds can be broadly categorized into two groups: unconditioned and conditioned. RGAN [2] is a classic unconditioned point cloud generation method, it uses a GAN structure with several MLPs to adversarially generate point clouds. WarpingGAN [58] is a lightweight and efficient network, which warps several uniform 3D grids into point clouds with various resolutions. ShapeGF [6] first introduces the score-based network to unconditionally generate point clouds, it takes noise as input and aims to learn the implicit surface of 3D shapes. Conditioned generation tasks usually generate point clouds from RGB images [14], text information [40] or partial point clouds [41, 69, 34]. Fan et al[14] proposed a PointOutNet to predict multiple plausible point clouds when conditioned on RGB images. Point-E [40] first synthesizes a single-view RGB image using a text-to-image diffusion model, and then uses another diffusion model to generate a point cloud based on the generated image. The VRCNet [41] generates point clouds conditioned on partial point clouds, it considers the completion task as a relational reasoning problem and learns to reason about the spatial relationships between the points and complete the missing parts accordingly. PVD [69] and PDR [34] are diffusion-based networks, they also operate under conditioned scenarios, taking partial point clouds as conditions to complete the missing parts.
Point-based Human Modeling. Ma et al[35] initially endeavored to represent clothed human body surface by learning point clouds, which can represent thin structures and open surfaces. POP [36] proposes a novel dense point cloud representation that can model high-quality clothing shapes with various styles. Lin et al[31] first learned implicit templates to represent coarse clothing topology and then generated point sets to enhance clothing deformations.
HaP also utilizes point clouds to represent the human body. It reimagines human body reconstruction by reframing classic implicit approaches as an explicit point cloud generation task. Utilizing depth maps and SMPL models inferred from RGB images, HaP adopts a diffusion-based framework to produce detailed and complete human body point clouds, offering a fresh perspective on human body reconstruction.
Figure 3: The pipeline of our framework HaP. HaP first estimates a depth map and an SMPL model from the RGB input, which serve as conditions of a diffusion process to generate sparse human point clouds. At the refinement stage, we further propose a refinement network $\texttt{PNet}_{\Theta_{2}}(\cdot)$ and a depth replacement operation to enhance the quality of $\mathcal{H}_{\mathrm{coarse}}$, and finally reconstruct meshes from $\mathcal{H}_{\mathrm{final}}$ via screened Poisson [25]. PC: Point Cloud. FPS: Farthest Point Sampling [46].
II-D Clothed Human Reconstruction
Implicit-based Approaches. PIFu [50] represents a pioneering effort in integrating implicit functions for human reconstruction. Building upon this, PIFuHD [51] improves performance on high-resolution images through a multi-level architecture, trained with both front and back normal maps. Arch [23] and Arch++[19] focus on learning implicit functions within canonical spaces, enabling the reconstruction of animatable human bodies. Additionally, Xiu et al.[65] and Zheng et al.[68] incorporate SMPL models to provide prior knowledge of human body structure. Alldieck et al.[3] enhance visual fidelity by computing albedo and shading details, introducing an innovative rendering loss function. S3Fs [12] enables the creation of relightable and animatable avatars.
Explicit-based approaches. SiCloPe [39] employs a generative adversarial network (GAN) conditioned on RGB images to synthesize consistent silhouettes and generate textured 3D meshes. More recently, Xiu et al.[64] introduced ECON, which simultaneously addresses both implicit and explicit reconstruction by generating 2.5D meshes from learned frontal and back normal maps. Tang et al.[57] designed a two-stage 3D convolutional network to explicitly learn the TSDF volumes of the human body.
Depth-guided approaches. Several methods [62, 63, 24, 7] have explored the role of depth information in 3D human body reconstruction. FAX [53] converts single-view images into depth and albedo maps in 2D space to produce full 3D scans. Gabeur et al[17] trained a GAN to predict both front and back depth maps, converted them into point clouds, and used the screened Poisson method to reconstruct a human body. Similarly, Han et al. [18] introduced a two-stage depth estimation network that learns front and back depth maps based on different body parts. Wang et al.[59] trained an adversarial network to denoise the depth collected by RGB-D cameras and generate geometric details of the back view. Xiong et al.[63] proposed a depth-guided self-supervised strategy to learn the implicit surface with signed distance function (SDF) values instead of occupancy values. Chan et al.[9] designed a depth-oriented sampling scheme to learn high-fidelity geometric details. They achieved this by incorporating depth values as supplementary features at the pixel level, which served as the input for implicit functions.
The proposed HaP introduces a novel explicit reconstruction framework by integrating depth information and the SMPL model. This approach utilizes a diffusion model to generate a complete human body point cloud within 3D space. Furthermore, the depth replacement technique enhances geometric details in the front view, bridging the gap between depth-guided and explicit reconstruction methods and improving performance over current approaches.
III Proposed Method
III-A Overview
The overall processing pipeline of the proposed single-view 3D human reconstruction framework consists of two sub-tasks: (1) inferring and rectifying 3D geometric information from the input 2D RGB image; and (2) generating a high-quality 3D human point cloud conditioned on the produced 3D geometric information, after which mesh reconstruction can be directly performed.
As illustrated in Fig. 3, our HaP begins with two parallel branches for depth and SMPL estimation, which show complementary characteristics in terms of inferring 3D information of human geometry. Specifically, depth estimation is able to capture detailed geometric patterns of the front depth map $\mathcal{I}_{\mathrm{depth}}$ but requires additional efforts to recover invisible body parts; differently, the estimated SMPL $\mathcal{S}$ provides strong human body priors and avoids non-human structures (e.g., unnatural pose, degenerate body, broken limb) but suffers from fixed-topology and over-smoothed surface details. Thus, we are motivated to jointly exploit such two sources of 3D information. Particularly, we propose an optimization-based SMPL rectification module to align the initially estimated SMPL model $\mathcal{S}$ with the depth-inferred partial 3D point cloud $\mathcal{P}$. After unifying the geometric information of depth and rectified SMPL $\widehat{\mathcal{S}}$ into a 3D point cloud, which could still be problematic due to the inevitable gap between depth and SMPL cues, we use a conditional point diffusion model, which takes both the depth-inferred sparse partial point cloud $\mathcal{P}_{\mathrm{sparse}}$ and SMPL-inferred sparse point cloud $\widehat{\mathcal{S}}_{\mathrm{sparse}}$ as inputs to produce a more accurate 3D point cloud of complete human body point cloud $\mathcal{H}_{\mathrm{coarse}}$. In addition, a refinement stage is explored to further enhance the geometric quality of $\mathcal{H}_{\mathrm{coarse}}$ and get $\mathcal{H}_{\mathrm{final}}$. Finally, we apply a classic surface reconstruction algorithm [25] to directly extract a mesh representation.
III-B Estimating 3D Information from Single 2D Images
Given a single-view 2D RGB image denoted as $\mathcal{I}$, we aim to extract two 3D geometric information from it, i.e., the depth information $\mathcal{I}_{\mathrm{depth}}$ and the SMPL model $\mathcal{S}$.
III-B1 Depth estimation
As a classic task in computer vision, monocular depth estimation [38] has been extensively investigated and achieved remarkable advancements, which provides a generic way of lifting 2D information captured by RGB images to spatial structures in the 3D geometric space.
In our implementation, we employ the recent state-of-the-art monocular depth estimator MIM [62] to predict the corresponding depth map $\mathcal{I}_{\mathrm{depth}}$ of $\mathcal{I}$. Then we project the depth value at each 2D-pixel coordinate $(u,v)$ to a 3D point $(x,y,z)$ according to the camera’s parameters. (This work adopts the setting of PIFu [50], assuming a fixed camera model with parameters to convert depth maps into 3D partial point clouds, without predicting camera parameters for individual images.) Besides, we also append the original RGB value of each pixel, producing a colored 3D point denoted as $\textbf{p}=(x,y,z,r,g,b)$. Thus, by leveraging rembg [1] (or methods like Segment Anything [26]) to obtain the valid mask of the human body, we are able to extract the useful pixels and subsequently eliminate the invalid ones, consequently obtain a 3D partial human body point cloud denoted as $\mathcal{P}=\{\textbf{p}\}$.
III-B2 SMPL Estimation and Rectification
To provide explicit human shape priors and supplement the missing body parts in the partial depth-inferred point cloud $\mathcal{P}$, we estimate from $\mathcal{I}$ an SMPL model $\mathcal{S}$ following [67, 16]. However, as analyzed in previous works [68], there would be inevitable misalignment between $\mathcal{S}$ and $\mathcal{P}$ caused by depth ambiguity, as shown in Fig. 4 (b), especially under the single-image setup, which can directly degrade the quality of the whole human reconstruction pipeline. To address this issue, assuming that depth estimation usually shows high accuracy and thus can deduce $\mathcal{P}$ with satisfactory quality [^1] In some extreme cases where the estimated depth map shows poor quality, one can switch to reliability-aware depth estimation models (e.g., [20, 61, 22]) and adaptively determine the utilization of depth information., we propose a novel SMPL rectification module to further promote the alignment of $\mathcal{S}$ with $\mathcal{P}$ by updating the SMPL parameters, the process is illustrated in Fig. 4 (c).
Let $\mathcal{S}_{v}=(\mathcal{V}_{v},~{}\mathcal{F}_{v})$ and $\overline{\mathcal{S}}_{v}=(\overline{\mathcal{V}}_{v},~{}\overline{\mathcal{F}}_{v})$ be the visible and the invisible parts of $\mathcal{S}$ (i.e. $\mathcal{S}=\mathcal{S}_{v}\cup\overline{\mathcal{S}}_{v}$), respectively, where $\mathcal{V}_{v}$, $\overline{\mathcal{V}}_{v}$, $\mathcal{F}_{v}$, and ¯$\overline{\mathcal{F}}_{v}$ stand for the vertex and face sets of the two parts, and $\widehat{\mathcal{S}}$ the rectified SMPL model. Generally, this problem can be achieved through iterative rectifications to the shape parameters $\bm{\beta}$ and pose parameters $\bm{\theta}$ of $\mathcal{S}$ under the objective of minimizing the distance between $\mathcal{S}_{v}$ and $\mathcal{P}$. Besides, to avoid any unwanted scenarios where the invisible part gets unnaturally close to $\mathcal{P}$, resulting in unnatural poses, we add a regularization term to maximize the distance between $\mathcal{P}$ and $\overline{\mathcal{S}}_{v}$. We also regularize shape parameters $\bm{\beta}$ due to their inherent zero-mean nature as established [5]. Moreover, the optimization is subject to two constraints: (1) the difference between the 2D keypoints of $\widehat{\mathcal{S}}$ and $\mathcal{S}$, preventing excessive variations in the pose of $\widehat{\mathcal{S}}$ on the 2D image, and (2) the silhouette difference between the mask of $\widehat{\mathcal{S}}$ (i.e., $\mathcal{M}_{\widehat{\mathcal{S}}}$) when projected onto the 2D image and the human mask from $\mathcal{I}$ (i.e., $\mathcal{M}_{h}$), preserving the overall human shape of $\widehat{\mathcal{S}}$. In all, we explicitly formulate the problem as
$\operatorname*{{\color[rgb]{0,0,0}min}}_{\bm{{\color[rgb]{0,0,0}\beta}},\bm{{\color[rgb]{0,0,0}\theta}}}~{}{\color[rgb]{0,0,0}\mathcal{L}_{\mathrm{r1}}}{\color[rgb]{0,0,1}=}$ ${\color[rgb]{0,0,0}\lambda_{1}\texttt{P2F}(\mathcal{P},~{}\mathcal{S}_{v})+\lambda_{2}\texttt{CD}(\mathcal{V}_{v},\mathcal{P})}-$ (1)
${\color[rgb]{0,0,0}\lambda_{3}\texttt{P2F}(\mathcal{P},\overline{\mathcal{S}}_{v})}{\color[rgb]{0,0,0}+\lambda_{4}\bm{\beta}+\lambda_{5}\texttt{L}_{1}(\mathcal{K},\mathcal{K}_{0})},$
[…]
Figure 4: Illustration of the SMPL rectification process. (a) Initially estimated SMPL model. (b) Registration before rectification. (c) Rectification process. (d) Registration after rectification. The directly estimated SMPL model is not well registered with the partial point cloud, and the situation is significantly relieved after the rectification process.
Figure 5: Illustration of the point cloud generation process. (a) Diffusion process. (b) Refinement process.
III-C Diffusion-based Explicit Generation of Human Body
The depth-inferred partial human body point cloud $\mathcal{P}$ offers free-form yet detailed geometric structures such as wrinkles in clothes of the visible side, leaving the occluded side completely absent. Instead, the rectified SMPL parametric human shape surface $\widehat{\mathcal{S}}$ is ensured to provide a complete body surface with reasonable pose and shape, while its fixed-topology causes essential difficulties in dealing with loose and complex clothing. These complementary characteristics motivate us to jointly exploit $\mathcal{P}$ and $\widehat{\mathcal{S}}$ for reconstructing high-fidelity 3D human geometry. Nevertheless, directly merging together the depth-inferred and the SMPL-inferred 3D geometric information can be problematic, since there usually exists a non-negligible gap between $\mathcal{P}$ and $\widehat{\mathcal{S}}$.
To achieve this, we propose training a conditional denoising diffusion probabilistic model (DDPM) to generate high-quality 3D human body point clouds. The model aims to naturally merge $\mathcal{P}$ and $\widehat{\mathcal{S}}$, eliminating any gaps. We sample sparse point clouds from $\mathcal{P}$ and $\widehat{\mathcal{S}}$ using farthest point sampling (FPS) [46], denoted as $\mathcal{P}_{\mathrm{sparse}}$ and $\widehat{\mathcal{S}}_{\mathrm{sparse}}$, respectively. The backbone of the diffusion model is a two-branch PointNet++ architecture [34, 46], parameterized by $\Theta_{1}$, denoted as $\texttt{PNet}{\Theta_{1}}(\cdot)$. As shown in Fig. 5 (a), a condition feature extractor extracts features from the sparse condition point clouds $\mathcal{P}_{\mathrm{sparse}}$ and $\widehat{\mathcal{S}}_{\mathrm{sparse}}$, while a time embedding layer extracts features of the diffusion step $t$. The features from the condition point clouds and the time step are then concatenated and used as input to the diffusion network. The diffusion network takes the noisy point cloud $\mathcal{H}_{\mathrm{gt}}^{t}$ and outputs the Gaussian noise $\epsilon_{t}$ at this diffusion step.
In our experiments, we found that although the diffusion stage was implemented, it did not completely remove all noise, resulting in overly rough geometric information that was not suitable for direct reconstruction. To address this, we introduce a refinement stage after the diffusion process to improve the quality of the generated point cloud by eliminating the remaining noise. As illustrated in Fig. 5 (b), we use a refinement network, parameterized by $\Theta_{2}$, denoted as $\texttt{PNet}{\Theta_{2}}(\cdot)$, to refine the generated human point clouds $\mathcal{H}_{\mathrm{coarse}}$ by learning displacement $\Delta$ for each point. Additionally, we apply a depth replacement strategy to densify the generated point cloud.
III-C1 Diffusion Stage
At this stage, we train a conditional diffusion model to generate a human point cloud. Due to constraints imposed by limited GPU memory and computational time, we configure $\texttt{PNet}_{\Theta_{1}}(\cdot)$ to generate a relatively sparse $\mathcal{H}_{\mathrm{coarse}}$ with 10,000 points. During the training process, we use the farthest point sampling operation to sample points from $\mathcal{P}$ and $\widehat{\mathcal{S}}$ respectively, which are concatenated as conditions for the diffusion model. A conditional DDPM consists of two Markov chains in opposite directions, i.e., the forward process and the reverse process, both processes have the same sampling step length $T$. During the forward process, the Gaussian noise is added to $\mathcal{H}_{\mathrm{gt}}$ at each sampling step $t$, and when $T$ is large enough, the output of the forward process should have a distribution similar to the Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$, the forward process is from the initial step $\mathcal{H}_{\mathrm{gt}}^{0}$ to the last step $\mathcal{H}_{\mathrm{gt}}^{T}$. Note that the condition $(\mathcal{P},\widehat{\mathcal{S}})$ is not included in the forward process.
The sampling process of the forward direction is iterative, however, we cannot train the DDPM at each time step due to the significant computational cost. According to [21], we can achieve arbitrary step $\mathcal{H}_{\mathrm{gt}}^{t}$ by defining $1-\gamma_{t}=\alpha_{t}$, $\bar{\alpha_{t}}=\prod_{i=1}^{t}\alpha_{i}$, where $\gamma$ is a pre-defined hyperparameter. Thus we formulate $q(\mathcal{H}_{\mathrm{gt}}^{t}|\mathcal{H}_{\mathrm{gt}}^{0})=\mathcal{N}(\mathcal{H}_{\mathrm{gt}}^{t};\sqrt{\bar{\alpha_{t}}}\mathcal{H}_{\mathrm{gt}}^{0},(1-\bar{\alpha_{t}})\bm{I})$, and $\mathcal{H}_{\mathrm{gt}}^{t}$ can be sampled through:
$\mathcal{H}_{\mathrm{gt}}^{t}=\sqrt{\bar{\alpha_{t}}}\mathcal{H}_{\mathrm{gt}}^{0}+\sqrt{1-\bar{\alpha_{t}}}\bm{\epsilon},$ (2)
where $\bm{\epsilon}$ is Gaussian noise. With a parameterization technique proposed in [21], we can train $\Theta_{1}$ by minimizing:
$\mathcal{L}_{\mathrm{ddpm}}=$ $\mathbb{E}_{t,\mathcal{H}_{\mathrm{gt}}^{t},\epsilon}\|\epsilon-\texttt{PNet}_{\Theta_{1}}(\mathcal{H}_{\mathrm{gt}}^{t},\mathcal{P},\widehat{\mathcal{S}},t)\|^{2}$ (3)
During the training process, both $\mathcal{P}$ and $\widehat{\mathcal{S}}$ are included as conditions at each time step to provide supervision on clothing information and human body information. At the inference process, we generate $\mathcal{H}_{\mathrm{coarse}}$ following the reverse process, i.e., progressively sampling $\mathcal{H}_{\mathrm{gt}}^{t-1}$ from $\mathcal{H}_{\mathrm{gt}}^{t}$ where $t$ starts from $T$. The generated $\mathcal{H}_{\mathrm{coarse}}$ possesses the same clothing information as the provided $\mathcal{P}$, and its pose remains consistent with the pose provided by $\widehat{\mathcal{S}}$.
III-C2 Refinement Stage
During inference, we observe that it is typically impossible for $\texttt{PNet}_{\Theta_{1}}(\cdot)$ to eliminate all the noise in $\mathcal{H}_{\mathrm{coarse}}$. Consequently, we can only recover a coarse surface from $\mathcal{H}_{\mathrm{coarse}}$. To overcome this problem, we train $\texttt{PNet}_{\Theta_{2}}(\cdot)$ parameterized with $\Theta_{2}$ to learn a displacement for each point in $\mathcal{H}_{\mathrm{coarse}}$, i.e., $\Delta=\texttt{PNet}_{\Theta_{2}}(\mathcal{H}_{\mathrm{coarse}},\mathcal{P},\widehat{\mathcal{S}})$, to further remove the remaining noise in $\mathcal{H}_{\mathrm{coarse}}$, which enhances the detailed geometric information. We obtain the human point cloud $\mathcal{H}$ with $\mathcal{H}=\mathcal{H}_{\mathrm{coarse}}+\Delta$. We formulate this problem as:
$\min_{\Theta_{2}}\texttt{D}(\mathcal{H},\mathcal{H}_{\mathrm{gt}})+\alpha\texttt{R}_{\mathrm{smooth}}(\Delta),$ (4)
where $\texttt{D}(\cdot,\cdot)$ stands for a distance metric realized by [49], $\texttt{R}_{\mathrm{smooth}}(\cdot)$ is a spatial smoothness regularization term written as:
$\texttt{R}_{\mathrm{smooth}}(\Delta)=\frac{1}{3N_{\mathrm{src}}K_{s}}\sum\limits_{x\in\mathcal{H}_{\mathrm{coarse}}}\sum\limits_{x^{\prime}\in\mathcal{N}(x)}\|\delta_{x}-\delta_{x^{\prime}}\|_{2}^{2},$ (5)
where $N_{\mathrm{src}}$ is the number of points contained in $\mathcal{H}_{\mathrm{coarse}}$, $\mathcal{N}(\cdot)$ returns $K_{s}$ nearest neighbor points, and $\delta_{x}\in\Delta$ stands for the displacement of a typical point.
Figure 6: Illustration of the pipeline of the proposed depth replacement strategy. We first find dense partial point cloud $\mathcal{P}$’s nearest points within refined human point cloud $\mathcal{H}$ as $s_{1}$, and then we iteratively find $s_{1}$’s nearest points within $\mathcal{P}$, starting with $s_{1}$, then $s_{2}$, finally we achieve the final partial point cloud $\mathcal{H}_{\mathrm{final}}$, which is formed by combining occluded point cloud $\mathcal{H}-s_{1}$ and visible point cloud $s_{3}$.
In addition, in the preceding diffusion stage, we sample a relatively sparse set of 3D points as inputs to the DDPM for saving memory and computational costs, which determines the sparsity of $\mathcal{H}_{\mathrm{coarse}}$, as likewise $\mathcal{H}$. To produce a dense point cloud with enhanced geometric details, we specialize a simple yet effective depth replacement strategy that fully leverages the density of the preceding depth-inferred dense partial point cloud $\mathcal{P}$. As illustrated in Fig. 6, our strategy commences by employing the ball query operation [46] to locate the nearest points of $\mathcal{P}$ within refined human point cloud $\mathcal{H}$; these nearest points are subsequently removed and denoted as $s_{1}$. Following this, we identify the nearest points of the occluded point cloud $\mathcal{H}-s_{1}$ within $\mathcal{P}$ and designate them as $s_{2}$. Finally, we determine the nearest points to $s_{2}$ within $\mathcal{P}$ and label them as visible point cloud $s_{3}$. The final human point cloud, denoted as $\mathcal{H}_{\mathrm{final}}$, is formed by combining occluded point cloud $\mathcal{H}-s_{1}$ and visible point cloud $s_{3}$.
After obtaining the final human point cloud $\mathcal{H}_{\mathrm{final}}$, we directly adopt Screened Poisson [25] for surface extraction.
IV Experiments
IV-A Data Preparation
One of the primary development bottlenecks in the research area of single-view human reconstruction is the issue of data insufficiency. To address this issue, we have created a new dataset, referred to as CityUHuman [^2] https://drive.google.com/drive/folders/1N-0UeC4_oU4R5lrn5avn_oZvaNp
m1uBG?usp=sharing
, which expands the number of available human body scans. We will make this dataset publicly available for research purposes so that it serves as a valuable supplement for assessing and benchmarking the performance of single-view human reconstruction methods. In comparison to the Thuman dataset [54, 66], the human models in the CityUHuman dataset exhibit smoother and more accurate surface details. As shown in Fig. 7, the Thuman2 scan displays a disrupted geometry at the pocket area due to its rough surface, whereas the CityUHuman scan preserves the finer geometric details more accurately.
[…]
Specifically, we obtained 3D human body scans using a high-precision, portable 3D scanner named Artec Eva [^3] https://www.artec3d.com/portable-3d-scanners/artec-eva. The complete collection of each scan comprises over 10 million points. We invited in total twelve volunteers and each volunteer was collected 3 to 11 scans. Adhering to ethical principles and respecting personal information ownership, we obtained consent from each volunteer prior to the data collection process. Volunteers were informed that their 3D-scanned data would be exclusively available for non-commercial research applications. In an effort to protect the privacy of the volunteers, we kindly request users to blur the faces in any materials intended for publication, such as papers, videos, and posters. Example scans with different clothes and poses from our dataset are illustrated in Fig. 2 to showcase its quality and potential for use in research.
IV-A1 Training data
In our experiments, we uniformly adopted Thuman2.0 [66] for the training of our proposed HaP learning architecture. On the one hand, we extracted from each model 10000 3D points as the ground-truth point cloud. On the other hand, we employed the Blender software for rendering ground-truth depth maps and photo-realistic images with an interval of 10 degrees. Thus, for each RGB image serving as the input, we have three aspects of supervision signals, i.e., a depth map, an SMPL-inferred point cloud, and a human point cloud. The total amount of the resulting training samples is 18000.
In particular, for making fair comparisons with IntegratedPIFu [9], we further created a subset from our training samples to train our HaP by preserving the same scan indices used in the training process of IntegratedPIFu, while evenly selecting 10 views from each scan.
IV-A2 Testing data
We generated testing data from Thuman3.0 [54] and our collected CityUHuman datasets to evaluate the performances of different approaches. We rendered each scan using 8 views spanning every 45 degrees in the yaw axis, and then randomly selected 459 RGB images from Thuman3.0 and 40 RGB images from CityUHuman. In addition, we also collected several real images from the Internet and inferred the corresponding human bodies for a purely qualitative evaluation.
[…]
IV-B Implementation Details
The employed depth estimation method MIM takes RGB images with a resolution of $512\times 512$ as inputs. During the training process of MIM, we employed the swin-v2-base architecture [32] as the backbone. The learning rate was set to $0.3\times 10^{-4}$, the batch size was set to 8, and the number of epochs was set to 30. The remaining parameters were kept the same as those in the original MIM. The training was performed on four RTX 3090 Ti GPUs.
During the training process of PNet, since the SMPL models lack texture information, we set their color values to 0, and there was no color information present in the ground truth human point clouds ($\mathcal{H}_{\mathrm{gt}}$). At the refinement stage, we generated 18000 coarse human point cloud samples $\mathcal{H}_{\mathrm{coarse}}$ and randomly selected 1600 of these to train the refinement network to save time. Both the diffusion stage and the refinement stage had a batch size and epoch of 8 and 300, respectively. The remaining parameters were kept the same as in the original PNet [34].
When rectifying the predicted SMPL models, we utilized the SGD optimizer to perform error backpropagation directly on the SMPL parameters $\bm{\beta},~{}\bm{\theta}$ for a total of 2000 iterations. The learning rate for the SGD was set to 0.03 and the momentum was set to 0.9.
We used FPS to uniformly sample 8192 points for $\mathcal{P}$, and 1024 and 2048 points for $\widehat{\mathcal{S}}$ at the diffusion stage and the refinement stage, respectively. We set the value of $k$ in the depth replacement module to 30 and the depth of Screened Poisson to 8.
Figure 12: Comparisons of the results refined (a) with only P2F loss, (b) without keypoint regularization and (c) with all regularizations.
TABLE I: Quantitative comparisons of different methods on two datasets. $\downarrow$ indicates that the lower values are better.
Method Metric Thuman3.0 CityUHuman
[…]
PIFu [50] 1.323 1.461 2.742 1.515 1.745 2.907
ICON [65] 1.306 1.395 2.731 1.315 1.333 2.813
ECON [64] 1.467 1.524 2.877 1.433 1.506 2.794
Our HaP 0.860 0.925 2.170 1.039 1.173 2.371
IntegratedPIFu [9] 1.148 1.182 2.262 1.140 1.187 2.306
Our HaP-I 1.021 1.065 2.468 1.154 1.194 2.523
2K2K [18] 1.405 1.384 8.184 1.258 1.304 8.052
Our HaP-T 0.827 0.887 2.140 1.043 1.174 2.371
TABLE II: Quantitative comparisons of visible and occluded sides on Thuman3.0. $\downarrow$” indicates that the lower values are better.
Method Metric Visible Side Occluded Side
[…]
PIFu [50] 1.381 1.376 1.993 1.421 1.463 2.347
ICON [65] 1.555 1.575 2.010 1.589 1.727 2.364
ECON [64] 1.634 1.702 1.884 1.618 1.755 2.283
Our HaP 0.960 1.070 1.450 1.043 1.206 1.780
IntegratedPIFu [9] 1.310 1.438 0.987 1.264 1.252 1.892
Our HaP-I 1.150 1.246 1.736 1.218 1.374 1.948
2K2K [18] 1.972 2.331 3.688 1.739 1.935 3.919
Our HaP-T 0.907 1.001 1.464 1.009 1.176 1.766
IV-C Comparisons with State-of-the-Art Methods
We primarily compared the proposed HaP with PIFu [50] [^4] Note that the PIFu method used in this comparison is implemented by Xiu et al. [65] for its better performance than the original implementation., ICON [65] and ECON [64]. Additionally, we compared HaP with IntegratedPIFu [9] and 2K2K [18], which are proposed for processing images of resolutions 1k and 2k, respectively. (All methods are trained on Thuman2.) Following Saito et al[50] and Xiu et al[65], we evaluated the performance of different methods using three metrics: CD (Chamfer Distance), P2F (Point-to-Face) distance, and normal difference. To compute CD, we sampled points uniformly on both the ground truth and reconstructed meshes and calculated the average bi-directional point-to-face distances. The P2F distance is calculated as the average distance from ground truth points to the reconstructed mesh. To compute the normal difference, we render the ground truth and reconstructed meshes from four fixed views using Pytorch3D, and calculate the L2 distances between them. This captures high-frequency geometric details.
Generally, as listed in Table I, our HaP demonstrates superior quantitative performance compared to PIFu [50], ICON [65], ECON [64], and 2K2K [18] on both Thuman3.0 and CityUHuman datasets by a large margin. And HaP can achieve competitive performance comparable to IntegratedPIFu [9]. Specifically, compared with existing fully-implicit learning approaches [50, 65, 9], as visualized in Fig. 8, HaP can correctly reconstruct poses and exhibits superior local geometric detail recovery. Additionally, Fig. 9 showcase some instances where PIFu [50] and ICON [65] encounter challenges, such as self-occlusion and loose clothing (marked with RED circle), while our HaP successfully addresses them.
Note that, IntegratedPIFu [9] typically demands high-resolution images of $1024\times 1024$ as input, resulting in extensive rendering times for preparing the training data of our HaP. To facilitate a quantitative comparison, we trained and tested HaP with the same scans but a lower spatial resolution, i.e., $512\times 512$. This method is referred to as HaP-I. Despite the data difference, our HaP-I still outperforms IntegratedPIFu in terms of CD and P2F on Thuman3.0, as reported in Table I.
In addition to fully-implicit learning, recent studies have also explored fully explicit and hybrid processing pipelines, such as ECON [64] and 2K2K [18]. Both of them aim to combine two parts of a human body, whereas HaP aims to generate a human body based on a partial point cloud. While ECON effectively addresses the issue of loose clothing, it still faces challenges related to its rendering-based SMPL refinement method. This can introduce SMPL inaccuracies when ECON combines two partial 2.5D meshes, resulting in unexpected outcomes, as illustrated in Fig. 9 (marked with RED circle). With our proposed SMPL rectification, HaP consistently outperforms ECON, both quantitatively and qualitatively. Moreover, 2K2K [18] usually fails to reconstruct human bodies when the Openpose [8] cannot detect all human keypoints due to self-occlusion. Therefore, omitting the failure cases of 2K2K, we conducted a re-evaluation of our HaP, denoted as HaP-T. The quantitative results listed in Table I indicate that our HaP-T outperforms 2K2K on both the Thuman3.0 and CityUHuman datasets. It is worth mentioning that the visual comparative results shown in Fig. 9 reveal the limb absence of the reconstructed surface at occlusion boundaries, We consider that 2D keypoints fail to provide accurate 3D human body structural information effectively. (marked with RED circle) Compared to 2K2K [18], our HaP-T can accurately reconstruct the entire human body while maintaining reasonable surface quality.
Moreover, we further extended the quantitative evaluation by separately measuring the visible and occluded sides of the reconstructed meshes in Table II, where HaP basically achieves large performance gains against all the competing methods, except for IntegratedPIFu on the Normal metric for the visible side as well as the P2F and Normal metrics on the occluded side. However, it is worth pointing out again that the input image for IntegratedPIFu is of higher resolution (i.e., $1024\times 1024$), while ours is of $512\times 512$.
Finally, to assess the effectiveness of our HaP in real-world scenarios, we gathered a collection of images featuring diverse clothing styles and poses from the internet to infer their full 3D structures. Fig. 10 visually compares the results of our HaP with ICON and ECON, where it can be seen that HaP can correctly reconstruct the clothing details and poses, while ICON cannot reconstruct the loose clothing details and ECON fails to recover correct human poses. More results of our HaP are depicted in Fig. 11, demonstrating commendable performance in realistic scenarios. We also refer readers to the video demo [^5] https://youtu.be/x8OtDLgkBHU.
TABLE III: Comparison of different SMPL rectification methods. Thuman2 is used for evaluation.
Depth Normal PaMIR only P2F w/o Keypoint Ours
CD $\downarrow$ 1.664 1.502 1.243 1.056 0.823 0.818
P2F $\downarrow$ 1.714 1.548 1.319 1.100 0.869 0.860
Normal $\downarrow$ 2.354 2.271 2.064 1.929 1.559 1.554
Figure 13: Visual comparisons of different SMPL rectification methods. (a) Initial predicted. (b) Depth-based rendering. (c) Normal-based rendering. (d) PaMIR. (e) Ours. (f) Ground Truth.
IV-D Analysis of SMPL Rectification
The accuracy of an SMPL model is pivotal for the precise reconstruction of the 3D human body. Nonetheless, as illustrated in Fig. 13, SMPL models directly predicted from RGB images frequently exhibit incorrect poses when viewed in 3D space. Xiu et al. [65] sought to tackle this problem by introducing a rendering-based optimization method. However, this approach still possesses certain limitations.
IV-D1 Defects of rendering-based optimization
Xiu et al. [65] employed a technique involving the rendering normal maps of SMPL models; subsequently, they attempted to optimize the SMPL models by minimizing the L1 distance between these SMPL normal maps and estimated normal maps of RGB images. However, this approach shows some inherent limitations. The pixels in normal maps are often not perfectly aligned, making the L1 distance an imperfect metric for quantifying the dissimilarities between the two normal maps. Consequently, the optimized SMPL models may exhibit inaccuracies and unnatural poses. Despite the incorporation of human prior information, ICON’s performance doesn’t exhibit significant improvements when compared to PIFu, as illustrated in Table I. This highlights the negative impact of imprecise SMPL models on the overall performance.
In addition to optimizing SMPL models with normal maps, we further experimented with depth maps. The qualitative and quantitative comparisons are presented in Fig. 13 and Table III, respectively.
Figure 14: The error maps and partial point clouds of depths predicted by (a) P3Depth [43] and (b) MIM [62].
Figure 15: The point clouds generated by (a) MIM+PDR and (b) MIM+VRCNet.
Figure 16: (a) Predicted partial point cloud. (b) Diffusion-based generated point cloud. (c) Point cloud after depth replacement.
IV-D2 The effectiveness of optimization in 3D space
Several methods [60, 4] have evaluated the efficacy of registering SMPL models within the implicit 3D space. In the context of monocular human reconstruction, PaMIR [68] also optimizes the SMPL models in the implicit 3D space by minimizing the occupancy loss to make the SMPL align with the predicted implicit function, as shown in Table III and Fig. 13, it achieves better quantitative and qualitative results than rendering-based methods.
[…]
Our proposed SMPL rectification method outperforms the rendering methods by accurately registering with the partial point cloud. This leads to better performance for HaP using our efficient and effective SMPL rectification method.
Figure 17: Visual comparison between the results by HaP (a) with and (b) without the SMPL rectification module. We present two views of the SMPL models and the reconstructed surfaces, respectively.
Figure 18: (a) The score boxplot of different methods (Green Triangle is the average scores, Red Line is the mid-value.) (b) Times of IntegratedPIFu and HaP being selected as the most realistic per person.
TABLE IV: The ablation studies on different modules of HaP. “SR” and “DR” are the SMPL rectification and depth replacement modules, respectively. Thuman3.0 is used for evaluation. The row highlighted in blue corresponds to our final HaP. $\downarrow$” means the lower, the better.
Depth Point SMPL SR Refinement Stage CD $\downarrow$ P2F $\downarrow$ Normal $\downarrow$
DR $\texttt{PNet}_{\Theta_{2}}$
P3Depth PNet PyMAF 0.946 1.015 2.301
MIM VRCNet PyMAF 1.643 1.275 9.352
MIM PNet PIXIE 0.883 0.944 2.203
MIM PNet PyMAF 1.051 1.079 2.622
MIM PNet PyMAF 0.877 0.944 7.082
MIM PNet PyMAF 0.958 0.978 7.437
MIM PNet PyMAF 0.860 0.925 2.170
Figure 19: Illustration of the limitation resulted from the flawed SMPL prediction. (a) Depth-inferred point cloud. (b) Refined SMPL model. (c) Generated point cloud. (d) Replaced point cloud. (e) Reconstructed surface.
IV-E Ablation Studies
We conducted comprehensive ablation studies on Thuman3.0 to gain a better understanding of the four modules encompassed in our HaP. For the depth estimation module, the conditioned diffusion-based generation module, and the SMPL estimation module, we considered three alternative methods: P3Depth [43], VRCNet [41], and PIXIE [16], respectively.
IV-E1 Depth estimation
Note that MIM [62] achieves more accurate depth on Thuman3.0 than P3Depth [43], i.e., 0.0158 vs. 0.0189 in terms of RMSE. From the $1^{st}$ row of Table IV, as expected, a better depth estimation method can lead to better reconstruction performance.
However, as visualized in Fig. 14, both of them contain many noisy points, which are further removed by our conditioned diffusion-based generation module.
IV-E2 Conditioned diffusion-based generation
We replaced PNet [34] with VRCNet [41], a point cloud completion network, to evaluate the effectiveness of the diffusion-based generation network.
As illustrated in Fig. 15, VRCNet generates highly noisy point clouds, which adversely affects the accuracy of the depth replacement module and leads to poor performance of HaP, as indicated in the $2^{nd}$ row of Table IV, with a higher normal error. Moreover, our diffusion-based point cloud generation is able to effectively remove the noisy points produced by depth estimation (see the noisy points between legs in Fig. 16).
IV-E3 SMPL estimation and rectification
Compared to PyMAF [67], PIXIE [16] (the $3^{rd}$ row) slightly under-performs in terms of CD and Normal metrics. As listed in the 4th row of Table IV, although removing the SMPL rectification module does not drastically reduce the numerical results, the head looks rather weird and unnatural, as exemplified in Fig. 17.
Moreover, we conducted experiments by replacing the SMPL models used in ICON and ECON with our HaP-rectified SMPL models. The modified methods are denoted as ICON-S and ECON-S. As shown in Table V, a comparison between ICON with ICON-S and ECON with ECON-S reveals that HaP-refined SMPL models can improve the performance of ICON and ECON, demonstrating the effectiveness of SMPL rectification. Besides, Table V shows that HaP consistently outperforms ICON-S and ECON-S, underscoring the effectiveness of our explicit diffusion-based framework over the implicit function models utilized by ICON and ECON.
IV-E4 Refinement stage
As listed in the $5^{th}$ row of Table IV, although HaP without the depth replacement strategy achieves satisfactory performance in terms of CD and P2F, there is a significant increase in the Normal error, demonstrating its lack of effectiveness in recovering surface details.
Furthermore, we assess the impact of omitting the displacement-based refinement module, $\texttt{PNet}{\Theta_{2}}$, which improves the point cloud’s geometric details by learning point-wise displacements. As shown in the $6^{th}$ and $7^{th}$ rows of Table IV, the substantial performance drop across all metrics when $\texttt{PNet}{\Theta_{2}}$ is removed highlights its essential role in producing high-quality results.
TABLE V: Quantitative comparisons of ICON, ICON-S, ECON, ECON-S, and our Hap on Thuman3 when the same SMPL pose parameters obtained by our HaP were adopted by all methods.
CD P2F Normal
ICON 1.306 1.395 2.731
ICON-S 0.982 1.054 2.484
ECON 1.467 1.524 2.877
ECON-S 0.970 1.034 2.447
HaP 0.860 0.925 2.170
IV-F Perceptual Study
We conducted two perceptual evaluations to quantitatively compare various methods. It’s important to note that 2K2K [18] was excluded from this evaluation, as it typically struggles to reconstruct the human body in cases where OpenPose [8] fails to detect all keypoints.
In the first perceptual evaluation, the questionnaire [^6] Perceptual Evaluation 1: https://forms.gle/23Hv4cDGbki56w7Z7 consists of ten questions. Each question presented an RGB image and six videos. The first video in each question displayed the ground truth scanned model, while the remaining five videos corresponded to five different methods. The order of the methods was randomized across the ten questions. Participants were asked to rate the methods on a scale from 1 to 5 based on their similarity to the real model. For instance, if method A was judged as the most accurate, it would receive a score of 5; if it was considered the least accurate, it would receive a score of 1. An intuitive guide for scoring was provided [^7] Intuitive Figure: https://forms.gle/M9X2EiyrJ7EfXp6d8. Participants were not filtered via catch trials, and no special training was provided. The evaluation focused on aspects such as posture accuracy, limb completeness, clothing style reconstruction, rationality of the non-visible side, and surface smoothness. As shown in Fig. 18 (a), HaP achieved the highest mean score.
In the second perceptual evaluation [^8] Perceptual Evaluation 2: https://forms.gle/kzybD19NronWAiec6, participants were shown the same videos and asked to select the best method between IntegratedPIFu and HaP, as they had obtained the highest scores in the first evaluation. Fig. 18 (b) shows that HaP was selected as the best method in six out of the ten samples.
IV-G Discussion on Limitations
Depth Estimation. While we assume that depth estimation methods generally produce accurate depth maps for human bodies, we acknowledge that these methods can occasionally fail. Since our approach relies on depth estimation, we recognize that suboptimal results may occur when depth estimation is inaccurate.
SMPL Rectification. HaP is also constrained by the possibility of imperfect SMPL predictions. While our SMPL rectification method offers improved efficiency compared to rendering-based techniques, there are situations where it may not effectively rectify the SMPL models, even when the estimated depth maps are accurate. This can lead to suboptimal reconstruction results. For example, as illustrated in Fig. 19, HaP faces challenges in aligning the SMPL model’s knee with the partial point cloud, resulting in a noticeable difference in the reconstruction outcome.
Smooth Occluded Side. Another limitation is the overly smooth geometry on the occluded side. Predicting occluded geometry is particularly challenging due to the lack of occluded information. In this study, we did not utilize occluded side normal or depth map predictions, leaving the occluded areas unsupervised, whereas other methods [64, 65, 9, 18] use. In future work, we plan to explore methods for incorporating appropriate supervision for the occluded side to achieve more realistic and natural geometry.
Loop-Closure Artifact. Some scans in the CityUHuman dataset may exhibit “loop-closure” artifacts. This occurs because, during the scanning process, the scanner had to be repositioned to complete the human body scan. This process, which typically lasts a few minutes, could result in small movements by the volunteer, causing repetitions in certain areas.
V Conclusion
We presented HaP, a novel point-based learning pipeline for single-view 3D human reconstruction. After coarsely inferring 3D information from the 2D image domain, our approach is featured by explicit point cloud manipulation, generation, and refinement in the original 3D geometric space. Compared with previous mainstream implicit field-based approaches, the proposed HaP is highly flexible, generalizable, robust, and capable of modeling fine-level geometric details. Extensive experiments demonstrate the superiority of our HaP over the current state-of-the-art methods. It is foreseeable that with the continuous development of various foundational modules and architectures such as depth estimation and diffusion models, the overall performance of our HaP pipeline will also improve progressively.